# Dynamic Indexing Through Learned Indices with Worst-case Guarantees

### v1

@@ Abstract
[…]
Since indexing queries are decomposable, the PGM index may be made dynamic through the logarithmic method. This method partitions $S$ into $O(\log n)$ buckets $B_{i}$ and stores a static PGM index $(h_{i},A_{i})$ for each bucket. It has amortised $O(\log n)$ update time and, insertion only, answers queries in $O(k+\sum\limits_{i}^{\lceil\log n\rceil}(\varepsilon+\log|h_{i}|))$ time. When allowing deletions, range query times deteriorate to worst-case $O(N+\sum\limits_{i}^{\lceil\log n\rceil}(\varepsilon+\log|h_{i}|))$ time (where $N$ is the largest size of $S$ seen so far).
This paper offers a combination of theoretical insights and experiments as we apply techniques from computational geometry to dynamically maintain a learned index with worst-case guarantees. First, we observe that testing whether there exists a learned index of $S$ that is a single segment reduces to deciding whether two convex hulls intersect. We combine this observation with the dynamic convex hull data structure by Overmars and van Leeuwen to dynamically maintain an approximately minimum-complexity learned index $h:\mathcal{U}\rightarrow[0,n]$ with $O(\log^{2}n)$ update time.
[…]
@@
Algorithms Engineering, indexing, data structures, convex hulls
@@ 1 Introduction
In the fundamental problem of data indexing, one seeks to store a large data set in order to efficiently answer queries such as rank, predecessor, and range queries. The input is a set $S$ of $n$ distinct integers from a universe $\mathcal{U}$. Denote by $\textsc{RANK}:S\rightarrow[n]$ the function that maps each $s\in S$ to its index in the sorted order. The goal is to store $S$ subject to indexing queries:
[…]
  • predecessor($q$) $=\max\{t\in S\mid t<q\}$                                           we allow $q\not\in S$
  • rank($q$) $=\textsc{RANK}(\textnormal{{predecessor}}(q))+1$                                         we allow $q\not\in S$
Additionally, we consider range queries, denoting $k$ as their output size:
  • range($q$, $t$) $=S\cap[q,t]$                                                              we allow $q,t\not\in S$
We study the dynamic indexing problem where $S$ is subject to insertions and deletions. We will consider update time, query time and space usage. To this end, we note that storing $S$ in an array uses $n$ space and focus on the additional space that a data structure may use.
Static indexing.
Solutions to the indexing problem can largely be grouped into two categories. The first category consists of tree structures over $S$ [1, 3, 27, 31]. Statically, this tree can be as simple as storing $S$ in an array $A$ in order. This approach therefore uses zero additional space but carries a computational cost. Through tree traversals, this approach can answer indexing queries in logarithmic time and range queries in $O(\log n+k)$ time.
The second category consists of map-based solutions [4, 18, 26]. These store $S$ in an array $A$ in sorted order and create a (hash) map $H:S\rightarrow[n]$ that maps each element in $S$ to $\textsc{RANK}(s)$. This gives constant-time membership, predecessor and rank queries if we restrict query input to elements of $S$ only. It supports range($q$, $t$) in $O(k)$ time whenever we restrict $q$ and $t$ to lie in $S$. Map based solutions use $O(n)$ additional space.
Recently, an additional category called learned indices was introduced [7, 13, 14, 17, 19]. Given some integer $\varepsilon$ (e.g., $\varepsilon=64$), a learned index is a function $h:\mathcal{U}\rightarrow[0,n]$ such that $h(q)\in[\textnormal{{rank}}(q)-\varepsilon,\textnormal{{rank}}(q)+\varepsilon]$. An indexing structure is the combination of $h$, a search structure over $h$, and an array $A$ that stores $S$ in sorted order. These approaches use machine learning to compute $H$ and index $A$ with $h(q)$ to answer queries. Ferragina and Vinciguerra [13] observe that such a function $h$ is of a geometric nature. Indeed, $S$ has a corresponding two-dimensional integer point set $F_{S}$, obtained by mapping each value $s\in S$ to the point $(\textsc{RANK}(s),s)$. Machine learning is essentially a form of curve fitting on $F_{S}$.
[…]
@@
  • the solution uses only $O(|h|)$ additional space, and
  • $O(\varepsilon+\log|h|)$ is, for an appropriate choice of $\varepsilon$, efficient in practice.
There exist benchmarking papers for the performance of learned indices  [12, 19, 31, 32].
Figure 1: (a) a set of $n$ values $S$. (b) $S$ corresponds to an $xy$-monotone point set $F_{S}$. (c) The PGM index computes a $y$-monotone set of segments that starts and ends with a vertical halfline. For any query value $q$, $h(q)$ approximates the rank of $q$.
Dynamic indexing.
When $S$ is dynamic, it is not efficient to keep $S$ in a sorted array. Thus, all previous approaches have to adapt (see also Table 1).
Tree-based approaches can instead be stored as a pointer structure. These can be dynamically maintained using tree rotations in $O(\log n)$ time per update, while maintaining the same query times. The cost of this adaption is $O(n)$ additional space.
Map-based approaches support dynamic membership queries in constant update time.
Learned indices support all indexing queries dynamically through the logarithmic method by Overmars [24]. This method stores each element of $S$ exactly once across at most $\lceil\log n\rceil$ buckets, where the bucket $B_{i}$ has size $2^{i}$. Each bucket $B_{i}$ is either empty or full. Each full bucket $B_{i}$ stores a subset of $S$ in an array $A_{i}$ in sorted order. They also create for $B_{i}$ a learned index $h_{i}:\mathcal{U}\rightarrow[0,2^{i}]$. Queries decompose across the buckets:
@@
This way, indexing queries require only an additional factor $O(\log n)$ time. Let the learned index $h$ have a construction time of $T(n)$. This data structure can be maintained insertion-only in amortised $O(T(n)\log n)$ time. An insertion inserts a new value into $B_{0}$. Let $j$ be the maximum integer such that all $B_{i}$ for $i\in[0,j-1]$ are full. This approach empties these buckets, fills $B_{j}$ in sorted order, and constructs $(A_{j},h_{j})$ in $O(T(2^{j}))$ time.
Whenever we delete some $s\in S$, this approach instead inserts a tombstone $s^{*}$, which is a special copy of $s$. If an insertion fills a new bucket $B_{j}$, it first iterates over all elements. If $B_{j}$ contains both $s$ and $s^{*}$, it removes both elements. It then constructs $(A_{j},h_{j})$ twice. Once on all ‘normal’ values, and once on all tombstones in $B_{j}$. This way, deletions take the same time as insertions do. Indexing queries can be answered by combining queries to both the normal and tombstone structures. E.g., $\textnormal{{rank}}(q)$ is the rank of $q$ in the ‘normal’ data structure minus the rank of $q$ in the tombstone structure. This approach has three downsides:
@@
  • Second, this approach uses $O(N)$ space where $N$ is the largest size of $S$ seen so far.
  • Finally, this approach does not support output-sensitive range queries – as there may be $O(N)$ values $s$ that (together with their tombstones $s^{*}$) lie in between a query pair $(q,t)$.
Technique Space Update Queries Indexing queries Ranges Source
Tree $O(n)$ $O(\log n)$ all $O(\log n)$ $+\,O(k)$ [3]
Map $O(n)$ $O(1)$ member $O(1)$ [26]
Log. PGM $O(N)$ $O(\log n)$ all $O(\sum\limits_{i}^{\lceil\log n\rceil}(\varepsilon+\log|h_{i}|))$ $+\,O(N)$ [13]
Our PGM $O(n)$ $O(\varepsilon+\log^{2}n)$ all $O(\varepsilon+\log|h|)$ $+\,O(k)$ Thm. B.11
Table 1: The techniques for dynamic indexing queries. $N$ denotes the maximum size of $S$ over all updates. Orange running times are amortised. Dynamic maps support only membership queries.
Main contribution.
Our first contribution is interesting from both a theoretical and applied perspective. We observe that existing techniques from computational geometry can be adapted to dynamically maintain a PGM index $h:\mathcal{U}\rightarrow[0,n]$, where $|h|\leq\frac{3}{2}\cdot|h^{\prime}|$ for all PGM indices $h^{\prime}$, in worst-case $O(\log^{2}n+\varepsilon)$ time.
Specifically, we extend and adapt the classical convex-hull intersection testing argument by Chazelle and Dobkin [5] to test in $O(\log n)$ time whether there exists a segment $\ell$ that lies within $L_{\infty}$-distance of all points in $F_{S}$ (see Figure 2). We use the above test together with the dynamic convex hull algorithm by Overmars and van Leeuwen [25] to dynamically maintain a learned index $h$ such that there exists no PGM index $h^{\prime}$ with $|h|>\frac{3}{2}|h^{\prime}|$.
We do some algorithms engineering for convex hulls. The classical convex hull maintenance [25] and intersection algorithms [6] rely upon geometric predicates that, in order to be accurately evaluated, require real-valued computations. Any straightforward implementation will either contain errors, or, require an exact computation kernel (which is very slow). We adapt the logic of these algorithms to rely upon orientation testing of triangles instead – which can be done more robustly and efficiently. We subsequently implement our algorithms.
[…]
We empirically compare the performance of our dynamic PGM index to the one in [13]. Our empirical comparison looks at two measures: Counting the number of lines, and measuring time and space spent. For the first measure, we outperform state-of-the-art, obtaining very few lines. For the time/space efficiency measures, as expected, our update time, space usage and time for rank queries is larger – primarily because [13] can store all data contiguously in an array. Yet, our solution offers a significant improvement in range queries and space usage whenever the update sequence contains many deletions.
@@ 2 Preliminaries
Let $\varepsilon$ be a positive integer. A PGM index of $S$ is defined as a $y$-monotone set of segments that together cover the $y$-axis. We regard $h$ as a map from $y$-coordinates to $x$-coordinates and require that for all $q\in\mathcal{U}$, $h(q)\in[\textnormal{{rank}}(q)-\varepsilon,\textnormal{{rank}}(q)+\varepsilon]$.
Ferragina and Vinciguerra [13, Lemma 1] wrongfully claim an $O(n)$-time algorithm to compute a minimum complexity PGM index $h$. They invoke a streaming algorithm by O’Rourke [22] for fitting straight lines through data ranges. We show that this algorithm outputs a PGM index $h$ such that there exists no PGM index $h^{\prime}$ with $|h|>\frac{3}{2}h^{\prime}$ (see Appendix C). Their algorithm restricts $S$ to contain no duplicates. We assume the same setting and compute something slightly different as we define an $\varepsilon$-cover instead:
@@
Let $f$ be an $\varepsilon$-cover and $Q$ be a horizontal line with height $q\in[\min S,\max S]$. Let $(s,t)$ be the segment in $f$ closest to $q$. Then $(line(a,b)\cap Q).x\in[\textnormal{{rank}}(q)-\varepsilon,\textnormal{{rank}}(q)+\varepsilon]$.
@@ 3 Testing whether a set can be $\varepsilon$-covered by a single segment
We consider the following subproblem: given $\varepsilon$, a set of integers $S$ with no duplicates, and the edges of $CH(F_{S})$ in a balanced binary tree, can we compute in $O(\log n)$ time whether an $\varepsilon$-cover $f$ of complexity $1$ exists and, if so, can we output $f$? That is, we want to compute a line $\ell$ of slope at least $1$ such that all points in $F_{S}$ lie within $L_{\infty}$-distance $\varepsilon$ of $\ell$.
Denote by $L$ (or $U$) the point set obtained by shifting each $p\in F_{S}$ downwards and rightwards by $\varepsilon$ and adding (,)($\infty$,$\infty$)( ∞ , - ∞ ) (or upwards and leftwards by $\varepsilon$, adding (,)($\infty$,$\infty$)( - ∞ , ∞ ))
@@
Let $\ell$ be a line with slope at least $1$. All points in $F_{S}$ lie within $L_{\infty}$-distance $\varepsilon$ of $\ell$ if and only if $\ell$ lies below all points in $U$ and above all points in $L$.
@@
A line is an $\varepsilon$-cover of $S$ if and only if it has slope at least $1$ and separates $CH(L)$ and $CH(U)$.
From $\textnormal{CH}(F_{S})$ we can extract $\textnormal{CH}(L)$ and $\textnormal{CH}(U)$ in logarithmic time. Chazelle and Dobkin [6, Section 4.2] remark that hull intersection testing, in the negative case, ‘can be modified’ to output a separating line. In our case, all segments of the convex hulls have slope at least $1$ and so this line will be an $\varepsilon$-cover. Thus, we have reduced our algorithmic problem to the well-studied convex hull intersection testing problem and we are seemingly done.
However, the history of convex hull intersection testing is complex. Both Chazelle and Dobkin [6] and Dobkin and Kirkpatrick [9] have a claim to the first $O(\log n)$-time algorithm. In 1987, Chazelle and Dobkin [5] revisit their algorithm, giving a more elaborate description. In 1990, Dobkin and Kirkpatrick [10] revisit their algorithm, showing a unified approach for $O(\log^{2}n)$-time polyhedron intersection testing. O’Rourke [23] notes that the algorithm in [10] is incorrect. He repairs the argument, and gives an $O(\log^{2}n)$-time $C$-implementation of polyhedron intersection testing. Dobkin and Souvaine [8] revisit the problem once more in 1991, and note that previous works lacked the necessary details for implementation. Their C-implementation is workable, but not robust. In 2015, Barba and Langerman [2] note that the community still lacks a comprehensive algorithm for polyhedron intersection testing. As a side effect, they present an alternative $O(\log n)$-time algorithm for convex hull intersection testing using polar transformations. A master thesis by Walther [29], under the supervision of Afshani and Brodal, implements [2] and [8]. The source is no longer available. This 35-year history shows that convex hull intersection testing is both complicated and error-prone. As a consequence, there exists no robust, modern $O(\log n)$-time implementation. As far as we are aware, there is no published algorithm for obtaining a separating line in the negative case.
@@
In Appendix A we provide an $O(\log n)$-time, robust implementation based on the algorithm in [5], restricted to convex hulls of edges with positive slope that contain (,)($\infty$,$\infty$)( ∞ , - ∞ ) and (,)($\infty$,$\infty$)( - ∞ , ∞ ). We prove the correctness of our algorithm. We note that even in this heavily restricted case, our algorithms and analysis are nontrivial. We then develop an algorithm for computing a separating line in the negative case, as we show:
@@
4 Dynamically maintaining a learned index
We dynamically maintain a learned index $h$ of $S$ by maintaining an $\varepsilon$-cover $f$ of $S$. We guarantee that there exists no $\varepsilon$-cover $f^{\prime}$ of $S$ with $|f|>\frac{3}{2}|f^{\prime}|$. By Observation 2, we obtain a learned index $h$. By Observation 2, there exists no PGM index $h$ where $|f|>\frac{3}{2}|h|$.
Theorem 4.1.
Proof 4.2.
The proof is illustrated by Figure 3. For any $s,t\in\mathbb{Z}$ with $s\leq t$, we say that $S[s,t]$ is blocked if there exists no $\varepsilon$-cover of $S[s,t]$ of size $1$. We maintain an $\varepsilon$-cover $f$ where for all consecutive intervals $[a,b],[c,d]\in\Lambda(f)$, $S[a,d]$ is blocked. Thereby, $|f|\leq\frac{3}{2}|f^{\prime}|$ for any $\varepsilon$-cover $f^{\prime}$ of $S$ (for completeness, we give a proof of this fact in Appendix C).
[…]
Let $([w,x],[y,z],[a,s],[s,s],[s,b],[c,d],[e,f])$ be consecutive intervals in $\Lambda(f)$ and denote $I=([y,z],[a,s],[s,s],[s,b],[c,d])$ (see Figure 3 (c) ). For each $(s,t)\in I$, we have access to $T(S[s,t])$. For any consecutive pair $([s,t],[q,r])$ in $I$, we may join the trees $T(S[s,t])$ and $T(S[q,r])$ in $O(\log^{2}n)$ time to obtain $T([s,r])$. We invoke $T([s,r])$.get_hull() and apply Theorem 3.4 to test in $O(\log^{2}n)$ total time whether $S[s,r]$ is blocked. If it is not, we replace $[s,t]$ and $[q,r]$ by $[s,r]$. Otherwise, we keep $T(S[s,r])$ and a segment that is an $\varepsilon$-cover of $S[s,r]$.

### v2

# A Dynamic Piecewise-linear Geometric Index with Worst-case Guarantees

@@ Abstract
Recently, various papers study a special case where the the input data behaves in an approximately piece-wise linear way. Given the sorted (rank,value) pairs, and given some constant $\varepsilon$, one wants to maintain a small number of axis-disjoint line-segments such that, for each rank, the value is within ±$\pm\varepsilon$ of the corresponding line-segment. Ferragina and Vinciguerra (VLDB 2020) observe that this geometric problem is useful for solving indexing problems, particularly when the number of line-segments is small compared to the size of the dataset.
We study the dynamic version of this geometric problem. In the dynamic setting, inserting or deleting just one data point may cause up to three line-segments to be merged, or one line-segment to be split at most three-way. To determine and compute this, we use techniques from dynamic maintenance of convex hulls, and provide new algorithms with worst-case guarantees, including an $O(\log n)$ algorithm to compute a separating line between two non-intersecting convex hulls – an operation previously missing from the literature.
We then use our fully-dynamic geometry-based subroutine in an indexing data structure, combining it with a natural hashing technique. The resulting indexing data structure has theoretically efficient worst-case guarantees in expectation. We compare its practical performance to the solution of Ferragina and Vinciguerra, which was shown to perform better in certain structured settings [Sun, Zhou, Li VLDB 2023]. Our empirical analysis shows that our solution supports more efficient range queries in the special case where the update sequence contains many deletions.
@@
Algorithms Engineering, Data Structures, Indexing, Convex Hulls
@@ 1 Introduction
We investigate the use of learned indices for the design of dynamic indexing data structures.
Indexing data structures.
An indexing data structure maintains a set $S$ of $n$ distinct integers from a universe $\mathcal{U}$. Let $\textsc{RANK}:S\rightarrow[n]$ be the function mapping each $s\in S$ to its index in the sorted order of $S$. The objective is to support the following indexing queries:
  • member($q$) returns true if $q\in S$.
  • predecessor($q$) returns $\max\{t\in S\mid t<q\}$. (We allow $q\notin S$.)
  • rank($q$) returns $\textsc{RANK}(\texttt{predecessor}(q))+1$. (We allow $q\notin S$.)
Additionally, we consider range queries, where $k$ denotes the output size:
  • range($q$, $t$) returns $S\cap[q,t]$. (We allow $q,t\notin S$.)
Static indexing data structures fall into three broad categories: Tree-based solutions store $S$ in a sorted array $A$, requiring no additional space but incurring logarithmic query costs. Tree traversals enable predecessor, rank, and member queries in $O(\log n)$ time and range queries in $O(\log n+k)$ time [2, 4, 35, 40]. Map-based solutions store $S$ in sorted order and maintain a hash map $H:S\rightarrow[n]$ mapping each element to its rank [5, 24, 34]. This enables constant-time support for member, predecessor, and rank queries if queries are restricted to elements of $S$, and $O(k)$ time for range queries when both endpoints lie in $S$. The additional space is $O(n)$. A third category use what are called learned indices, a recently introduced term [25, 18, 17, 23, 9]. Given an integer parameter $\varepsilon$, a learned index is a function $h_{\varepsilon}:\mathcal{U}\rightarrow[0,n]$ such that
$h_{\varepsilon}(q)\in[\texttt{rank}(q)-\varepsilon,\texttt{rank}(q)+\varepsilon].$
The function $h_{\varepsilon}$ is learned from $S$ and used to guide search in a sorted array $A$ storing $S$. Ferragina and Vinciguerra [17] interpret $h_{\varepsilon}$ geometrically: each $s\in S$ maps to a point $(\textsc{RANK}(s),s)$ in the plane, and $h_{\varepsilon}$ is learned as a piecewise-linear approximation to $F_{S}$.
A notable instance of learned indices is the PGM index [17], where $h_{\varepsilon}$ is a $y$-monotone piecewise-linear function made of segments, with the property that each point in $F_{S}$ lies within an $\varepsilon$-wide horizontal strip around some segment. Let $|h_{\varepsilon}|$ denote the number of segments. The data structure supports indexing queries in $O(\varepsilon+\log|h_{\varepsilon}|)$ time and range queries in $O(\varepsilon+k+\log|h_{\varepsilon}|)$ time. They also show how to construct a PGM index in linear time, such that there exists no PGM index $h_{\varepsilon}^{\prime}$ with $|h_{\varepsilon}|>\frac{3}{2}|h_{\varepsilon}^{\prime}|$. Ferragina and Vinciguerra argue that learned indices are the ‘best of both worlds’ since:
@@
  • the solution uses only $O(|h_{\varepsilon}|)$ additional space, and
  • $O(\varepsilon+\log|h_{\varepsilon}|)$ is, for an appropriate choice of $\varepsilon$, efficient in practice.
Their performance has been empirically benchmarked in several studies [16, 25, 41, 40].
Figure 1: (a) a set of $n$ values $S$. (b) $S$ corresponds to an $xy$-monotone point set $F_{S}$. (c) The PGM index computes a $y$-monotone set of segments that starts and ends with a vertical halfline.
Dynamic indexing data structures.
Due to their fundamental role, dynamic indexing structures have received extensive theoretical and practical attention. When $S$ is dynamic, maintaining a sorted array becomes inefficient. Tree-based structures can be updated in $O(\log n)$ time with tree rotations. Map-based approaches allow constant-time member updates but are typically not extended to support other indexing queries. Learned indices offer a promising direction by exploiting structural properties of $S$, akin to parametrised algorithms. However, just as parametrised algorithms, data structures based on learned indices are not always efficient: if $S$ lacks exploitable structure or access patterns are skewed, traditional indexing data structures are preferred [36, 29]. Experimental studies have examined properties of learned indices [15, 14, 29, 36], in an effort to classify when they are appropriate to use. As a brief summary, learned indices pay a price in updates [29], and traditional indices are preferred if $S$ or the access pattern is complex or skewed, or if concurrency is possible [36]. If there is sufficient structure, both space usage and access times can benefit from learned indices [14, 15], subject to the strategy employed by the learned index.
Dynamic learned indices through the logarithmic method.
The logarithmic method of Overmars [32] provides an amortised way to maintain learned indices. $S$ is partitioned into $\lceil\log n\rceil$ buckets $B_{i}$, each of size $2^{i}$. Each bucket is either full or empty, and stores its contents in an array $A_{i}$ in sorted order and maintains a learned index $h_{\varepsilon}^{i}$ over $A_{i}$.
Let the learned index $h_{\varepsilon}$ have a construction time of $T(n)$. This data structure can be maintained insertion-only in amortised $O(T(n)\log n)$ time. An insertion inserts a new value into $B_{0}$. Let $j$ be the maximum integer such that all $B_{i}$ for $i\in[0,j-1]$ are full. This approach empties these buckets, fills $B_{j}$ in sorted order, and constructs $(A_{j},h_{\varepsilon}^{j})$ in $O(T(2^{j}))$ time. Whenever we delete some $s\in S$, this approach instead inserts a tombstone $s^{*}$, which is a special copy of $s$. If an insertion fills a new bucket $B_{j}$, it first iterates over all elements. If $B_{j}$ contains both $s$ and $s^{*}$, it removes both elements. It then constructs $(A_{j},h_{\varepsilon}^{j})$ twice. Once on all ‘normal’ values, and once on all tombstones in $B_{j}$. This way, deletions take the same time as insertions do. In this paper, we consider the following open question:
“Can a learned index be dynamically be maintained with worst-case guarantees?”
Intermezzo: computing a line cover.
The interpretation of learned indices by Ferragina and Vinciguerra [17] translates to a geometric problem where the goal is to (approximately) cover a monotone set of two-dimensional points by a set of line segments. Using the logarithmic method, and the static algorithm of O’Rourke [30] to approximately points, they dynamically maintain an $\varepsilon$-cover: a set of lines that are guaranteed to be within an $\varepsilon$ horizontal distance from each point. We consider the problem of maintaining dynamic $\varepsilon$-covers to be an interesting geometric problem in its own right.
From learned indices to indexing data structures.
Under the logarithmic method, indexing queries decompose naturally across the buckets:
@@
This way, indexing queries require only an additional factor $O(\log n)$ time. Indexing queries can be answered by combining queries to both the normal and tombstone structures. E.g., $\textnormal{{rank}}(q)$ is the rank of $q$ in the ‘normal’ data structure minus the rank of $q$ in the tombstone structure. This approach has two downsides:
@@
  • Second, this approach does not support output-sensitive range queries – as there may be $O(n)$ values $s$ that (together with their tombstones $s^{*}$) lie in between a query pair $(q,t)$.
This leads to the following open question:
“Can a dynamic learned index be converted into an output-sensitive dynamic indexing data structure with worst-case guarantees?”
Contribution and organization.
We propose maintaining a dynamic $\varepsilon$-cover (and thereby a dynamic learned index) via dynamic convex hull techniques. Section 3 shows that deciding whether $S$ admits an $\varepsilon$-cover of complexity 1 is equivalent to convex hull intersection testing (Figure 2). Section 4 shows a robust algorithm to compute the intersection between two convex hulls in $O(\log n)$ time. We adapt our algorithm to output a separating line (which is a learned index of complexity $1$) in the negative case. Section 5 combines these with dynamic convex hull data structures to yield a dynamic learned index worst-case $O(\log^{2}n)$ update time. We empirically compare our learned index to the PGM index from [17].
Section 6 introduces a novel hashing-based approach to convert learned indices into dynamic indexing data structures, using $O(\varepsilon^{-1})$ additional expected overhead. We compare our dynamic indexing structure to the amortised PGM index of [17] in terms of update time and index complexity. We do not benchmark against traditional indexing data structures – since the relation between learned and traditional indices is previously studied [36]. Instead, our goal is to push the theoretical limits and worst-case guarantees of learned indices.
@@ 2 Preliminaries
Let $\varepsilon$ be a positive integer. A PGM index $h_{\varepsilon}$ of $S$ is defined as a $y$-monotone set of segments that together cover the $y$-axis. We regard $h_{\varepsilon}$ as a map from $y$-coordinates to $x$-coordinates and require that for all $q\in\mathcal{U}$, $h_{\varepsilon}(q)\in[\textnormal{{rank}}(q)-\varepsilon,\textnormal{{rank}}(q)+\varepsilon]$.
Ferragina and Vinciguerra [17, Lemma 1] wrongfully claim an $O(n)$-time algorithm to compute a minimum complexity PGM index $h_{\varepsilon}$. They invoke a streaming algorithm by O’Rourke [30] for fitting straight lines through data ranges. We show that this algorithm outputs a PGM index $h_{\varepsilon}$ such that there exists no PGM index $h_{\varepsilon}^{\prime}$ with $|h_{\varepsilon}|>\frac{3}{2}h_{\varepsilon}^{\prime}$ (see Appendix C). Their algorithm restricts $S$ to contain no duplicates. We assume the same setting and compute something slightly different as we define an $\varepsilon$-cover instead:
@@
Let $f$ be an $\varepsilon$-cover and $Q$ be a horizontal line with height $q\in[\min S,\max S]$. Let $(s,t)$ be the segment in $f$ closest to $q$. Then $(line(s,t)\cap Q).x\in[\textnormal{{rank}}(q)-2\varepsilon,\textnormal{{rank}}(q)+2\varepsilon]$.
@@ 3 Testing whether a set can be $\varepsilon$-covered by a single segment
We consider the following subproblem: given a parameter $\varepsilon$, a set $S$ of $n$ distinct integers, and the edges of $\textnormal{CH}(F_{S})$ stored in a balanced binary tree, can we compute in $O(\log n)$ time whether there exists an $\varepsilon$-cover $f$ of complexity 1? Formally, we seek a line $\ell$ of slope at least 1 such that all points in $F_{S}$ lie within $L_{\infty}$-distance $\varepsilon$ of $\ell$. Let $L$ (resp. $U$) denote the set obtained by shifting each $p\in F_{S}$ downwards and rightwards (resp. upwards and leftwards) by $\varepsilon$, and adding the point (,)($\infty$,$\infty$)( ∞ , - ∞ ) (resp. (,)($\infty$,$\infty$)( - ∞ , ∞ )).
@@
Let $\ell$ be a line of slope at least 1. Then all points in $F_{S}$ lie within $L_{\infty}$-distance $\varepsilon$ of $\ell$ if and only if $\ell$ lies below all points in $U$ and above all points in $L$.
@@
$\ell$ is an $\varepsilon$-cover of $S$ iff it has a slope $\geq 1$ and separates $CH(L)$ from $CH(U)$.
Given $\textnormal{CH}(F_{S})$, we can extract $\textnormal{CH}(L)$ and $\textnormal{CH}(U)$ in $O(\log n)$ time. Chazelle and Dobkin [7, Section 4.2] remark that, in the negative case, convex hull intersection testing can be modified to produce a separating line. In our setting, the hulls consist of segments with slope at least 1, and any such separator corresponds to an $\varepsilon$-cover. Thus, our problem reduces to the classical convex hull intersection problem and we are seemingly done.
However, the history of convex hull intersection testing is long and intricate. Both Chazelle and Dobkin [7] and Dobkin and Kirkpatrick [11] independently proposed the first $O(\log n)$-time algorithms. In 1987, Chazelle and Dobkin [6] presented a more detailed description of their method. Dobkin and Kirkpatrick revisited their own work in 1990 [12], proposing a unified $O(\log^{2}n)$-time algorithm for polyhedron intersection, which O’Rourke later identified as incorrect [31]. He corrected the argument and provided a $C$-implementation. Further work by Dobkin and Souvaine [10] noted that earlier implementations lacked robustness. More recently, Barba and Langerman [3] observed that the community still lacked a complete, robust algorithm for polyhedral intersection. They proposed an alternative $O(\log n)$ algorithm based on polar transformations. Walther’s master’s thesis [38], supervised by Afshani and Brodal, implemented both this and earlier methods, but the source code is no longer available.
This 35-year history highlights the complexity and subtlety of convex hull intersection testing. Despite its history, no robust and modern $O(\log n)$-time implementation is available. Moreover, no published algorithm explicitly computes a separating line in the negative case.
@@
In Appendix A, we present a robust $O(\log n)$-time algorithm for convex hull intersection testing. Our algorithm is specialised to convex hulls composed of positively sloped segments and including the points (,)($\infty$,$\infty$)( ∞ , - ∞ ) and (,)($\infty$,$\infty$)( - ∞ , ∞ ). We formally prove its correctness and adapt it to compute a separating line in the negative case, thereby constructing an $\varepsilon$-cover of complexity 1 when it exists. The later adaption and its analysis are nontrivial, and arguably (partly) fill a gap in the existing literature on convex hull intersection testing.